Two wires of same length are placed parallel to each other at a distance of 0.65 metres from each other.
Calculate the ratio of the force they exert on each other assuming 10 amperes current in the first wire and 5 amperes in the second wire
1
Expert's answer
2011-06-22T10:04:43-0400
Denote the current in the first wire as I1 (10 amperes) and the current in the second wire as I2 (5 amperes).

The force that act at the unit length is

Fu=μ0/4π (2I1 I2)/r=10-7 *2*10*5 / 0.65=153.84*10-7 N.